OpenStudy (anonymous):

Evaluate a4 for a = -3.

OpenStudy (ankit042):

you mean a^4 right?

OpenStudy (anonymous):

Yes

OpenStudy (ankit042):

this is simple a^2 = a*a a^3 =a*a*a a^4=a*a*a*a Now just substitute the value of a

OpenStudy (anonymous):

its a4 a=-3 so u plug that in for a (-3)4=-12 hope this helps! (=

OpenStudy (ankit042):

@caitlynmcdaniel question is actually a^4 not a4 she made a typo while typing..I have confirmed

OpenStudy (anonymous):

a^4 would be 4a i thnk. idk how i explained it was right if the problem was typed how i thought ankit042 your right